### v1

# Milnor number of invariant singularity: generalization of Chulkov’s inequality

Abstract.
In [3] S.P. Chulkov has proven that if the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $\mu(f)\geq p-1$ for any function germ invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ for prime pp. Here we present a generalization of this inequality for actions of arbitrary abelian groups.
1. Introduction
In [1], M. E. Kazarian has posed a following conjecture [^1] This is essentially a local version of the Matsumura-Monsky theorem([5]). Actually the proof by Matsumura and Monsky can be easily adapted to this statement. However, Kazarian and others did not know that, and this is a good thing, as otherwise we would not have this interesting problem on Milnor numbers to wrestle with.:
Theorem $1.$
Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin and a compact Lie group $G$ of analytic diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ such that $f(g(x))=f(x)\;\forall\;g\in G$. If the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $G$ is discrete.
V.A. Vassiliev has given an elementary proof of this statement based on the geometry of the Newton polytope of a germ with an infinite group of symmetries in [2]. Some time later, S.P. Chulkov has given a proof ([3]) based on the following inequality for the Milnor number of $f$:
Theorem 1.2.
Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin such that the Taylor series of $f$ contains no monomials of degree $\leq 2$. If f is invariant with respect to a nontrivial action of a group of prime order $p$ then $\mu(f)\geq p-1$.
Kazarian’s conjecture follows from this easily, as an infinite compact Lie group contains cyclic subgroups of any order.
A slightly stronger inequality ($\mu(f)\geq p+1$) can be easily proven in case of a real group action using invariant morsifications as described in [4]. These two inequalities are essentially the only known bounds on the Milnor number in terms of properties of the group action, so it is interesting to obtain a bound for the Milnor number in case of groups of composite order. Chulkov’s original proof does not generalize (as far as we know) to groups other that $\mathbb{Z}_{p}$.
The main theorem of this paper is the generalization of Chulkov’s inequality to arbitrary abelian groups. For a group action $\tau$ on $(\mathbb{C}^{n},0)$ we will denote the length of shortest orbit of a non-fixed point by $l(\tau)$.
Theorem 1.3.
Let $f$ be a germ of analytic function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin such that the Taylor series of $f$ contains no monomials of degree $\leq 2$. Let $\tau$ be a nontrivial action of a finite abelian group $G$ on $(\mathbb{C}^{n},0)$ by analytic diffeomorphisms. If $f$ is invariant with respect to $\tau$ then $\mu(f)\geq l(\tau)-1$.
In the case $G=\mathbb{Z}_{p}$ this reduces to the original Chulkov’s inequality.
2. Definitions and auxiliary statements
In a suitable local system of coordinates any finite group action on $(\mathbb{C}^{n},0)$ can be linearized (see [6]). So we wil consider all group actions to be induced by a linear representation of $G$. Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x$. $\lambda$ is called the weight of $f$. Invariant functions are equivariant with weight 1, of course.
For a function germ $f$ $J_{f}$ will denote the Jacobian ideal $\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{n}}>$. $Q_{f}$ is the Jacobian algebra $\mathbb{C}{x_{1},\dots,x_{n}}/J_{f}$. $Q_{f}$ always admits a monomial basis, i.e. the basis consisting of monomials
If $f$ is invariant with respect to an action of a group $G$, then $G$ also acts on $Q_{f}$. The representation of $G$ on $Q_{f}$ is called the equvariant Milnor number of $f$ and denoted $\mu_{G}(f)$. $\nu(f)$ is the multiplicity of the trivial representation in $\mu_{G}(f)$, i.e. the dimension of the space of invariant elements of $Q_{f}$. This subspace of invariant elements is denoted $Q^{G}_{f}$.
A group action on $(\mathbb{C}^{n},0)$ on is called real if it admits an invariant quadratic form of rank $n$.
The next two theorems are taken from [4].
Theorem 2.1.
If $f$ is invariant with respect to a real action of a finite group $G$ then there is a $G$-invariant deformation $f_{\lambda}$ of $f$ with only nondegenerate critical points.
The set $C$ of critical points of any invariant deformation $f_{\lambda}$ is $G$-invariant. The action of $G$ on $(\mathbb{C}^{n},0)$ induces a representation $\theta$ of $G$ on the space of $\mathbb{C}$-valued functions on $C$.
Theorem $2.$
If the deformation $f_{\lambda}$ has only nondegenerate critical points then the representations $\theta$ and $\mu_{G}(f)$ are isomorphic.
The last statement mean in particular that $\nu(f)$ is equal to the number of orbits of critical points in a invariant Morse approximation of $f$.
Theorem 2.3.
If the main theorem is true for group action with no fixed points outside of origin, it is true for all group actions.
Proof.
Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point and the origin.
[…]
Since $\nu(f)$ is equal to the number of orbits of critical points of an invariant Morse approximation, in the case of a real representation $\tau$ with no fixed points outside of origin we obtain the inequality $\mu(f)\geq(\nu(f)-1)l(\tau)+1$. We will refer to this formula as Roberts’ inequality.
3. Proof of main theorem
For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$. $\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$. If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$. It is easy to see that if $f$ has an isolated singularity at the origin then so does $f\oplus f$, and $\mu(f\oplus f)=(\mu(f))^{2}$. Furthermore, if $f$ is invariant with respect to an action $\tau$ then $f\oplus f$ is invariant with respect to the action $\tau_{\mathbb{R}}$.
Theorem 3.1.
If $p_{1},\dots,p_{\mu}$ form a monomial basis for the Jacobian algebra of $f$ then $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ form a monomial basis in the Jacobian algebra of $f\oplus f$.
Proof.
Start with an arbitrary monomial basis $q_{1}(x,y),\dots,q_{\mu^{2}}(x,y)$} for $Q_{f\oplus f}$. Each $q_{i}$ factors as as a product of monomials in $x$ and $y$: $q_{i}=q_{i1}(x)q_{i2}(y)$. For each $i$ $q_{i1}\notin J_{f}$, as otherwise $q_{i}\in J_{f\oplus f}$. Hence we have a decomposition $q_{i1}=\Sigma_{j}c_{ij}p_{j}(x)$. Similarly, each $q_{i2}$ is a linear combination of $p_{j}(y)$. Therefore $p_{i}(x)p_{j}(y)$ form a system of generators for $Q_{f\oplus f}$. As there are exactly $\mu^{2}$ of them, they form a basis.
Theorem 3.2.
$\nu(f\oplus f)\geq\mu(f)$.
Proof.
Consider a monomial basis $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ of $Q_{f\oplus f}$ as described in the previous theorem. Each $p_{i}(x)$ is an equivariant function with respect to the action $\tau$ with some weight $\lambda$. As $G$ acts by conjugate characters on $y$, $p_{i}(y)$ is equivariant with the inverse weight $\bar{\lambda}$. Hence  $\forall\;1\leq i\leq\mu\;p_{i}(x)p_{i}(y)\in Q^{G}_{f\oplus f}$.
Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$: $\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
As we now know $\mu(f\oplus f)=\mu^{2}(f)$, $\nu(f\oplus f)\geq\mu(f)$ and, obviously, $l(\tau_{\mathbb{R}})=l(\tau)$. Substituting all of these into Roberts’ inequalit gives us the inequality $\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$.
Solving this quadratic inequality for $\mu(f)$, we obtain $\mu(f)=1$(which is impossible given that the 2-jet of $f$ is zero) or $\mu(f)\geq l(\tau)-1$.

### v3

@@
-Milnor number of invariant singularity: generalization of Chulkov’s inequality
+Milnor number of an invariant singularity: generalization of Chulkov’s inequality
 Abstract.
-In [3] S.P.
-Chulkov has proven that if the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $\mu(f)\geq p-1$ for any function germ invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ for prime pp.
-Here we present a generalization of this inequality for actions of arbitrary abelian groups.
+S.P.
+Chulkov has proven that the Milnor number of a function germ $f$ with zero 2-jet invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ is at least p1p-1 for prime $p$.
+In this paper we prove a generalization of this inequality for actions of arbitrary abelian groups using M.
+Roberts’ results on invariant morsifications for invariants of real group actions.
 1. Introduction
@@
 However, Kazarian and others did not know that, and this is a good thing, as otherwise we would not have this interesting problem on Milnor numbers to wrestle with.:
-Theorem $1.$
-Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin and a compact Lie group $G$ of analytic diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ such that $f(g(x))=f(x)\;\forall\;g\in G$.
+Theorem 1.1.
+Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin and a compact Lie group $G$ of analytic diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ such that $f(g(x))=f(x)\;\forall\;g\in G\;\forall\;x\in\mathbb{C}^{n}$.
 If the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $G$ is discrete.
@@
 Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
-A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x$.
+A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x\;\forall\;g\;\in\;G$.
 $\lambda$ is called the weight of $f$.
@@
 the basis consisting of monomials
-If $f$ is invariant with respect to an action of a group $G$, then $G$ also acts on $Q_{f}$.
+If $f$ is invariant with respect to an action of a group $G$, the Jacobian ideal is invariant and therefore $G$ also acts on $Q_{f}$ by changes of coordinates: $g\circ h(x)=h(g^{-1}(x))$.
 The representation of $G$ on $Q_{f}$ is called the equvariant Milnor number of $f$ and denoted $\mu_{G}(f)$.
@@
 The set $C$ of critical points of any invariant deformation $f_{\lambda}$ is $G$-invariant.
-The action of $G$ on $(\mathbb{C}^{n},0)$ induces a representation $\theta$ of $G$ on the space of $\mathbb{C}$-valued functions on $C$.
-Theorem $2.$
+The action of $G$ on $(\mathbb{C}^{n},0)$ induces a representation $\theta$ of $G$ on the space of $\mathbb{C}$-valued functions on $C$ ($\theta(g)\circ h(x)=h(g^{-1}(x))$.
+Theorem 2.2.
 If the deformation $f_{\lambda}$ has only nondegenerate critical points then the representations $\theta$ and $\mu_{G}(f)$ are isomorphic.
@@
 Proof.
-Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point and the origin.
+Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
 […]
@@
 For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$.
-$\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$.
-If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
+$\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$. If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
 For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$.
@@
 Hence  $\forall\;1\leq i\leq\mu\;p_{i}(x)p_{i}(y)\in Q^{G}_{f\oplus f}$.
-Now we can conclude the proof of the main theorem.
-As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$: $\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
-As we now know $\mu(f\oplus f)=\mu^{2}(f)$, $\nu(f\oplus f)\geq\mu(f)$ and, obviously, $l(\tau_{\mathbb{R}})=l(\tau)$.
-Substituting all of these into Roberts’ inequalit gives us the inequality $\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$.
+Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
+$\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
+As we now know $\mu(f\oplus f)=\mu^{2}(f)$, $\nu(f\oplus f)\geq\mu(f)$ and, obviously, $l(\tau_{\mathbb{R}})=l(\tau)$. Substituting all of these into Roberts’ inequality gives us the inequality
+$\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$.
 Solving this quadratic inequality for $\mu(f)$, we obtain $\mu(f)=1$(which is impossible given that the 2-jet of $f$ is zero) or $\mu(f)\geq l(\tau)-1$.
+4. Tightness of the bound
+The next step is to produce examples showing the tightness of this bound.
+The trivial example is the 1-dimensional representation: if a character $\chi$ of group $G$ has order $k$ and $G$ acts on $\mathbb{C}$ via this character, the invariant function $x^{k}$ has the lowest possible Milnor number k1k-1.
+However, this is unsatisfactory, as we would ideally like to prove that a)the bound is tight in arbitrarily large dimensions; b)the Milnor number is bounded by the length of the shortest orbit when there are orbits of different lengths.
+To show this, we present the following example.
+Let $n$ be an odd positive integer, $g$ a generator of a group $\mathbb{Z}_{2^{n}+1}$, $h$ a generator of $\mathbb{Z}_{3}$, $\epsilon,\sigma$ – primitive roots of unity of order $2^{n}+1,3$, respectively .
+$\mathbb{Z}_{2^{n}+1}\times\mathbb{Z}_{3}$ acts on $\mathbb{C}^{n}$ in the following manner: $g\circ(x_{1},\dots,x_{n})$
+$(\epsilon^{(-2)}x_{1},\epsilon^{(4)}x_{2},\dots,\epsilon^{(-2)^{n-1}}x_{n-1},\epsilon x_{n})$, $h\circ(x_{1},\dots,x_{n})=\sigma(x_{1},\dots,x_{n})$.
+The polynomial $x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+\dots+x_{n-1}^{2}x_{n}+x_{n}^{2}x_{1}$ is invariant and has Milnor number $2^{n}$, the length of the shortest orbit (e.g.
+of the point $(0,0,\dots,0,1)$) is $2^{n}+1$, however the longest orbit (e.g.
+of $(1,1,\dots,1)$) has length $3(2^{n}+1)$

### v4

@@
 Abstract.
-S.P.
-Chulkov has proven that the Milnor number of a function germ $f$ with zero 2-jet invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ is at least p1p-1 for prime $p$.
-In this paper we prove a generalization of this inequality for actions of arbitrary abelian groups using M.
-Roberts’ results on invariant morsifications for invariants of real group actions.
+We prove a lower bound for the Milnor number of function germ invariant with respect to a finite abelian group action.
+It is shown that this bound is tight for functions of arbitrarily many variables.
+We also prove the function germs that reach this lower bound are equivariantly stable, i.e.
+invariant analogues of Morse singularities.
 1. Introduction
@@
 So we wil consider all group actions to be induced by a linear representation of $G$.
-Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
+Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\ldots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\ldots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
 A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x\;\forall\;g\;\in\;G$.
@@
 Invariant functions are equivariant with weight 1, of course.
-For a function germ $f$ $J_{f}$ will denote the Jacobian ideal $\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{n}}>$.
-$Q_{f}$ is the Jacobian algebra $\mathbb{C}{x_{1},\dots,x_{n}}/J_{f}$.
+For a function germ $f$ $J_{f}$ will denote the Jacobian ideal $\frac{\partial f}{\partial x_{1}},\ldots,\frac{\partial f}{\partial x_{n}}>$.
+$Q_{f}$ is the Jacobian algebra $\mathbb{C}\{x_{1},\ldots,x_{n}\}/J_{f}$.
 $Q_{f}$ always admits a monomial basis, i.e.
@@
 This subspace of invariant elements is denoted $Q^{G}_{f}$.
-A group action on $(\mathbb{C}^{n},0)$ on is called real if it admits an invariant quadratic form of rank $n$.
+A group action on $(\mathbb{C}^{n},0)$ is called real if it admits an invariant quadratic form of rank $n$.
 The next two theorems are taken from [4].
@@
 Proof.
-Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
+Consider the action $\tau:g\circ(x_{1},\ldots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\ldots,x_{k},\chi_{k+1}(g)x_{k+1},\ldots,\chi_{n}(g)x_{n}),\chi_{k+1},\ldots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
 […]
@@
 3. Proof of main theorem
-For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$.
+For a group action $\tau:g\circ(x_{1},\ldots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\ldots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\ldots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\ldots,\bar{\chi}_{n}(g)y_{n})$.
 $\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$. If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
-For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$.
+For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\ldots,x_{n})+f(y_{1},\ldots,y_{n})$.
 It is easy to see that if $f$ has an isolated singularity at the origin then so does $f\oplus f$, and $\mu(f\oplus f)=(\mu(f))^{2}$.
@@
 Theorem 3.1.
-If $p_{1},\dots,p_{\mu}$ form a monomial basis for the Jacobian algebra of $f$ then $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ form a monomial basis in the Jacobian algebra of $f\oplus f$.
+If $p_{1},\ldots,p_{\mu}$ form a monomial basis for the Jacobian algebra of $f$ then $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ form a monomial basis in the Jacobian algebra of $f\oplus f$.
 Proof.
-Start with an arbitrary monomial basis $q_{1}(x,y),\dots,q_{\mu^{2}}(x,y)$} for $Q_{f\oplus f}$.
+Start with an arbitrary monomial basis $q_{1}(x,y),\ldots,q_{\mu^{2}}(x,y)$} for $Q_{f\oplus f}$.
 Each $q_{i}$ factors as as a product of monomials in $x$ and $y$: $q_{i}=q_{i1}(x)q_{i2}(y)$.
@@
 Theorem 3.2.
-$\nu(f\oplus f)\geq\mu(f)$.
+$\nu(f\oplus f)=k_{1}^{2}+\ldots+k_{m}^{2}$, with $k_{i}$ being the multiplicities of different one-dimensional representations in the irreducible decomposition of $Q_{f}$.
 Proof.
-Consider a monomial basis $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ of $Q_{f\oplus f}$ as described in the previous theorem.
-Each $p_{i}(x)$ is an equivariant function with respect to the action $\tau$ with some weight $\lambda$.
-As $G$ acts by conjugate characters on $y$, $p_{i}(y)$ is equivariant with the inverse weight $\bar{\lambda}$.
-Hence  $\forall\;1\leq i\leq\mu\;p_{i}(x)p_{i}(y)\in Q^{G}_{f\oplus f}$.
-Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
+As proven in the previous theorem $Q_{f\oplus f}$ is isomorphic ( as a vector space) to the tensor product of $Q_{f}$ with itself.
+On the second factor of this tensor product (i.e.
+on monomials in $y_{i}$) $G$ acts by conjugate characters, so the action on the group on the second factor can be identified with the action on the conjugate space $Q^{*}_{f}$.
+According to a well-known result in representation theory (see, e.g., [7], pp.
+433–434) for each finite-dimensional representation $V$ the representation $V\otimes V^{*}$is naturally isomorphic to $Hom(V,V)$, and the invariant subspace of $V\otimes V^{*}$ is isomorphic to the space of equivariant endomorphisms of $V$.
+The dimension of the space of equivariant endomorphisms is $k^{2}_{1}+\ldots+k^{2}_{m}$ with $k_{i}$ being the multiplicities of irreducible representations in the decomposition of $V$.
+Since $\mu(f)=k_{1}+\ldots+k_{m}$, we have, in particular, $\nu(f\oplus f)\geq\mu(f)$.
+Now we can conclude the proof of the main theorem.
+As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
 $\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
@@
 The trivial example is the 1-dimensional representation: if a character $\chi$ of group $G$ has order $k$ and $G$ acts on $\mathbb{C}$ via this character, the invariant function $x^{k}$ has the lowest possible Milnor number k1k-1.
-However, this is unsatisfactory, as we would ideally like to prove that a)the bound is tight in arbitrarily large dimensions; b)the Milnor number is bounded by the length of the shortest orbit when there are orbits of different lengths.
-To show this, we present the following example.
-Let $n$ be an odd positive integer, $g$ a generator of a group $\mathbb{Z}_{2^{n}+1}$, $h$ a generator of $\mathbb{Z}_{3}$, $\epsilon,\sigma$ – primitive roots of unity of order $2^{n}+1,3$, respectively .
-$\mathbb{Z}_{2^{n}+1}\times\mathbb{Z}_{3}$ acts on $\mathbb{C}^{n}$ in the following manner: $g\circ(x_{1},\dots,x_{n})$
-$(\epsilon^{(-2)}x_{1},\epsilon^{(4)}x_{2},\dots,\epsilon^{(-2)^{n-1}}x_{n-1},\epsilon x_{n})$, $h\circ(x_{1},\dots,x_{n})=\sigma(x_{1},\dots,x_{n})$.
-The polynomial $x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+\dots+x_{n-1}^{2}x_{n}+x_{n}^{2}x_{1}$ is invariant and has Milnor number $2^{n}$, the length of the shortest orbit (e.g.
-of the point $(0,0,\dots,0,1)$) is $2^{n}+1$, however the longest orbit (e.g.
-of $(1,1,\dots,1)$) has length $3(2^{n}+1)$
+However, this is unsatisfactory, as we would ideally like to prove that the bound is tight in arbitrarily large dimensions.
+To show this, we present the following examples.
+Let $n$ be an odd positive integer, $d_{1},\ldots,d_{n}\geq 2$ – natural numbers, $g$ – generator of a group $\mathbb{Z}_{d_{1}\ldots d_{n}+1}$, $\epsilon$ – pimitive root of unity $1$ of degree $d_{1}\ldots d_{n}+1$.
+$\mathbb{Z}_{d_{1}\ldots d_{n}+1}$ acts on $\mathbb{C}^{n}$ in the following fashion:
+$g\circ(x_{1},\ldots,x_{n})=(\epsilon x_{1},\epsilon^{-d_{1}}x_{2},\ldots,\epsilon^{(-1)^{n-2}d_{1}\ldots d_{n-2}}x_{n-1},\epsilon^{(-1)^{n-1}d_{1}\ldots d_{n-1}}x_{n})$
+The polynomial $x_{1}^{d_{1}}x_{2}+x_{2}^{d_{2}}x_{3}+\ldots+x_{n-1}^{d_{n_{1}}}x_{n}+x_{n}^{d_{n}}x_{1}$ is invariant with respect to this action.
+It has an isolated critical point, as it is an invertible polynomial of the loop type (see., e.g., [8, 9]), and its Milnor number is equal to $d_{1}\ldots d_{n}$.
+The length of any nontrivial orbit in this case is $d_{1}\ldots d_{n}+1$, so the bound $\mu\geq l(\tau)-1$ is tight.
+5. Equivariantly stable germs
+There are many papers on the classification of equivariantly simple singularities (see., e.g., [10], [11], [12]).
+Roughly speaking, an equivariantly stable singularity is the simplest of the equivariantly simple singularities, i.e.
+the one that is only adjacent to itself.
+If the case of real group action the equivariantly stable singularity is the nondegenerate one, as the deformation of a Morse function is again a Morse function.
+For the formal definition of equivariant stability see below.
+We will denote $\mathfrak{m}_{G}$ the maximal ideal in the ring of $G$-invariant analytic germs.
+Since our action has no fixed points outside of origin, every element of $\mathfrak{m}_{G}$ has a critical point at the origin.
+Consider $\mathcal{O}^{r}_{G}$ — the space of $r$-jets of elements in $\mathfrak{m}_{G}$.
+The group $D^{r}_{G}$ of $r$-jets of $G$-equivariant diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ acts on this space.
+A germ $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is equivariantly stable if for all large enough $r$ $D^{r}_{G}$-orbit of $r$-jet of $f$ in $\mathcal{O}^{r}_{G}$ is open.
+Theorem 5.1.
+$f$ is equivariantly stable iff $\nu(f)=1$.
+Proof.
+Let $f$ be equivariantly stable.
+As the orbit of $r$-jet of $f$ is open for large enough $r$ the tangent space to the orbit has dimension equal to the dimension of $\mathcal{O}^{r}_{G}$ itself.
+The tangent space to the orbit at $r$-jet of $f$ is the projection of the jacobian ideal of $f$ to $\mathcal{O}^{r}_{G}$.
+But if for all $r$ large enough the projection of $J_{f}$ covers all $\mathcal{O}^{r}_{G}$ $J_{f}$ contains $\mathfrak{m}_{G}$.
+Since $f$ has a critical point at the origin $J_{f}$ does not contain 1 and $\nu(f)=1$.
+Now assume $\nu(f)=1$.
+The Milnor number of $f$ is finite, as the invariants of a finite group action have the separating property, i.e.
+for any two orbits of the $G$-action there is an invariant function that is equal to 1 on the first orbit and 0 on the second.
+But if $\nu(f)=1$, then $J_{f}=\mathfrak{m}_{G}$ and any invariant function takes the same value at any point of the singular locus of $f$ as it does at the origin.
+So the separating property fails if the critical point of $f$ is non-isolated.
+Since $f$ has a finite Milnor number we can apply the Slodowy theorem (see [13]), according to which the invariant versal deformation of $f$ is equal to $f+\lambda_{1}h_{1}+\ldots+\lambda_{\nu}h_{\nu}$, $h_{j}$ — linearly independent invariant elements of $Q_{f}$.
+Since $\nu(f)=1$, the versal deformation of $f$ is trivial, i.e.
+has the form $f_{\lambda}=f+\lambda$, and the equivariant stability follows immediately.
+∎
+In particular for the nondegenerate singularity $\mu(f)=\nu(f)=1$, as $Q_{f}$ is equal to $\mathbb{C}$.
+Theorem 5.2.
+If $\mu(f)=l(\tau)-1$ $f$ is equivariantly stable
+Proof.
+If $\mu(f)=l(\tau)-1$, the inequality $\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$ becomes an equality.
+Since also $\mu^{2}(f)\geq(\nu(f\oplus f)-1)l(\tau)+1$ and $\nu(f\oplus f)\geq\mu(f)$, it can only become an equality if $\nu(f\oplus f)=\mu(f)$.
+As was proven earlier, $\nu(f\oplus f)=k_{1}^{2}+\ldots+k_{m}^{2}$, with $k_{i}$ being the multiplicities of one-dimensional representations in $Q_{f}$.
+With $\mu(f)$ equal to $k_{1}+\ldots+k_{m}$, the equality $\nu(f\oplus f)=\mu(f)$ can only hold if $\forall\;i\;k_{i}=1$.
+In particular, the multiplicity of the trivial representation (equal to $\nu(f)$) is 1.
+∎